OpenStudy (anonymous):

\[\sum_{n=1}^{3}i^(n-1)\]

OpenStudy (anonymous):

I'm going to assume the n is representing natural numbers. Then it will be i(1-1)+i(2-1)+i(3-1)=0+i+2i=3i.